OpenStudy (anonymous):

multiply (x^6-x^4)(x^6+x^4)

OpenStudy (anonymous):

use \[(a+b)(a-b)=a^2-b^2\] with \[a=x^6,b=x^4\]

OpenStudy (anonymous):

x^12-x^8=x^4